OpenStudy (anonymous):

Help please Calculate the mass of each product formed when 236g Silver sulfide reacts with excess hydrochloric acid Ag2S + 2HCl = AgCl + H2S I already balanced it.

OpenStudy (anonymous):

The eqn is Ag2S + 2HCl = 2AgCl + H2S Mole Ag2S=mass/Mr=236/(108*2)+32 = 0.952 mole Mole HCl= 0.952*2=1.904 moles Mass HCl= Mole * Mr= 1.904* 36.5=69.496 g Mass AgCl= 1.904*(108+35.5)=273.224 g Mass H2S= 0.952*(2+32)=32.368 g